### v3

# Autoencoding-Free Context Compression for LLMs via Contextual Semantic Anchors

Abstract
Context compression presents a promising approach for accelerating large language model (LLM) inference by compressing long contexts into compact representations. Current context compression methods predominantly rely on autoencoding tasks to train context-agnostic compression tokens to compress contextual semantics. While autoencoding tasks enable compression tokens to acquire compression capabilities, compression via autoencoding tasks creates a fundamental mismatch: the models are optimized for reconstruction that diverge from actual downstream tasks, thereby weakening the features more beneficial for real-world usage. We propose Semantic-Anchor Compression (SAC), a novel method that shifts from autoencoding task based compression to an architecture that is equipped with this compression capability a priori. Instead of training models to compress contexts through autoencoding tasks, SAC directly selects so-called anchor tokens from the original context and aggregates contextual information into their key-value (KV) representations. By deriving representations directly from the contextual tokens, SAC eliminates the need for autoencoding training. To ensure compression performance while directly leveraging anchor tokens, SAC incorporates two key designs: (1) anchor embeddings that enable the compressor to identify critical tokens, and (2) bidirectional attention modification that allows anchor tokens to capture information from the entire context. Experimental results demonstrate that SAC consistently outperforms existing context compression methods across various compression ratios. On out-of-distribution evaluation using MRQA, SAC achieves 1 EM improvement at 5x compression over strong baselines, with increasing advantages at higher compression ratios.
++footnotetext: These authors contributed equally to this work.
1 Introduction
The expanding scope of large language models (LLMs) to tasks like processing long documents (Liu et al., 2024b; Li et al., 2024; Duan et al., 2025), maintaining multi-turn dialogue coherence (Zhang et al., 2025; Yi et al., 2025; Guan et al., 2025), and generating responses grounded in extensive external knowledge (Lewis et al., 2020; Karpukhin et al., 2020; Huang et al., 2025) necessitates the incorporation of vast contexts into the model input. However, directly processing such extremely long contexts is fraught with challenges, including prohibitive computational costs, significant inference latency, and performance degradation, largely caused by the “lost-in-the-middle” phenomenon (Liu et al., 2024a).
To address these challenges, recent studies have proposed context compression (Chang et al., 2024; Li et al., 2025a), a technique that typically appends special tokens (i.e. compression tokens) to the end of the context and leverages the LLM’s causal attention mechanism to compress contextual information into a compact representation within these tokens. Once this compact representation is obtained, the LLM can generate responses conditioned on it, rather than being conditioned on the entire original context. This significant reduction in context length leads to substantial decreases in both inference time and GPU memory consumption. While effective, these approaches face a key limitation: the compression tokens are randomly initialized and lack inherent semantic information. To compensate, they typically rely on extensive pretraining on both autoencoding (AE) and language modeling (LM) tasks (illustrated in Figure 1) to endow the compression tokens with the ability to carry contextual information. While AE task has shown to be necessary in ICAE (Ge et al., 2024) since compression tokens lack context-relevant semantics, the AE task requires the compressed representation to reconstruct all tokens in the context, even low-information tokens. This reliance on a suboptimal and costly pretraining stage raises a critical research question: Is it possible to design a compression architecture that inherently understands context without a demanding AE phase?
Figure 1: Three tasks for training the context compressor introduced by ICAE and followed by numerous works. The training uses (a) Autoencoding task and (b) Language modeling task to pretrain the encoder, then finetunes on (c) Question answering task.
To answer this question, this work introduces Semantic-Anchor Compression (SAC) (Figure 2), a novel architecture for the context compression task. Instead of appending new special tokens and requiring extensive autoencoding pretraining to learn their representations, SAC directly selects representative tokens from the original context to act as ‘anchor tokens’ for compression. By leveraging these semantically meaningful anchors from the input itself, SAC incorporates natural semantic priors that obviate the need for autoencoding pretraining. To signify their special role, these selected tokens are then augmented with dedicated ‘anchor embeddings’, enabling the LLM to distinguish them from regular tokens. Furthermore, to enhance their compression capabilities, we modify the standard causal attention to a bidirectional attention mechanism. This allows anchor tokens to access information from the entire context, rather than being restricted to only preceding tokens. These modifications collectively foster a more effective context compression by providing anchor tokens with both distinct representations and comprehensive contextual awareness. Empirically, we test SAC on the MRQA (Fisch et al., 2019a) dataset and confirm that it outperforms existing strong context compression baselines. For example, compared to 500xCompressor (Li et al., 2025b) at 5x compression, the average exact match (EM) improves from 25.4 to 32.3. Results show that 1) our proposed method improves more in absolute accuracy over strong baselines on more challenging high compress ratio scenarios 2) our proposed architecture achieves its best performance in a simpler training setting without autoencoding training arguably because the anchor tokens already contain enough information about the original context. Our analysis reveals that SAC’s compressed representations more closely resemble original context token KVs in feature space, so that LLMs performing inferenece can arguably better understand them.
Figure 2: The difference between SAC and previous methods. While previous methods (a) compress contextual information into context-agnostic special tokens (referred to as compression tokens), SAC (b) compresses the context directly into the original contextual tokens themselves. Here, $\tilde{M}$ can represent either the output from the final layer of the LLM or the Key-Value pairs, being later on used as compressed representations for LLM inference.
2 Related Works
2.1 Compression method
Many methods focus on reducing prompt lengths. CC (Wingate et al., 2022) utilizes contrastive learning to compress specific natural language prompts into shorter and unique soft prompt tokens. However, it cannot generalize to unseen prompts and requires retraining for new prompts. GIST (Mu et al., 2023) compresses original prompts into KV values through finetuning and can handle arbitrary unseen contexts. AutoCompressor (Chevalier et al., 2023) recursively combines compressed vectors with sub-prompts and aggregates all compressed vectors to construct the final representation, enabling compression of longer contexts. However, both GIST and AutoCompressor require finetuning the LLMs performing inference (referred to later as target LLM), which may affect LLMs’ original capabilities.
ICAE (Ge et al., 2024) formulates context compression as training a general encoder that compresses contexts into compact representations understandable by target LLMs without finetuning. To achieve this, ICAE introduces autoencoding tasks and performs joint pretraining with language modeling tasks, followed by finetuning on downstream tasks. 500xCompressor (Li et al., 2025b) improves upon ICAE by replacing the compression carrier from the last layer output of compression tokens with KV values at each layer, achieving higher compression ratios. EPL (Zhao et al., 2025) identifies that ICAE and 500xCompressor neglect the impact of positional encoding and proposes distributing compression token position IDs uniformly across the entire context rather than placing them at the end. However, these methods still rely on autoencoding tasks to endow the compression tokens with the ability to carry contextual information.
Another category of prompt compression methods is based on token selection, which selects representative tokens from contexts based on token importance. SelectiveContext (Li et al., 2023), LLMLingua (Jiang et al., 2023), and LongLLMLingua (Jiang et al., 2024) employ causal small language models to evaluate token importance based on information entropy. LLMLingua-2 (Pan et al., 2024a) distills a token classifier to compute the probability of each token to be preserved. These works demonstrate that LLMs can understand original contexts using a small number of representative tokens. However, they do not perform compressed tokens training which limits the usability of the selected tokens by target LLMs. Our proposed SAC can be seen as a combination of token selection methods and compressed token training methods: it derives and train compressed representations that are based on tokens selected directly from the context and indeed is compatible with the token selection methods above.
2.2 Bidirectional Attention
Recent studies have shown that, removing the decoder’s unidirectional causal constraint and introducing bidirectional attention can effectively enhance the model’s representational capacity (Wang et al., 2020). For instance, NV-Embed (Lee et al., 2025) replaces causal attention with bidirectional attention during contrastive training, achieving strong performance on general text embedding and dense vector retrieval tasks. LLM2Vec (BehnamGhader et al., 2024), by enabling bidirectional attention alongside masked next-token prediction, significantly improves the model’s ability to capture global semantics in text embedding tasks. These works indicate that bidirectional attention is advantageous for acquiring global semantic information and robust contextual representations. However, its effectiveness in text compression tasks remains underexplored. Motivated by these findings, we incorporate bidirectional attention into the compressor to enhance contextual modeling during the compression phase.
3 Method
3.1 Task Formulation
Context compression is formally defined as follows: an encoder $\mathcal{E}$ compresses a context $C=(c_{1},c_{2},\ldots,c_{|C|})$ into a compact representation $\tilde{M}$ with $\tilde{M}=\mathcal{E}(C)$. Subsequently, a target LLM leverages the compressed representation $\tilde{M}$ in place of the original context $C$ to perform various tasks, such as question answering.
[…]
Additionally, during finetuning, the question answering loss $\mathcal{L}_{\mathbf{QA}}$ enhances the ability of $\tilde{M}$ to extract information that is potentially relevant for downstream tasks (e.g. QA). Since the encoder operates without knowledge of what questions might be asked later, it learns to identify and preserve information that is likely to be queried, enabling accurate answer generation $A=(a_{1},a_{2},\ldots,a_{|A|})$ when presented with subsequent questions $Q=(q_{1},q_{2},\ldots,q_{|Q|})$, as shown in Figure 1c; mathematically, $\mathcal{L}_{\mathbf{QA}}=-\log P(A|\tilde{M},Q)$.
3.2 Semantic-Anchor Compressor
A key distinction between our approach SAC and previous methods is that we derive compressed representations directly from selected context tokens, as shown in Figure 2. This involves selecting a subset of tokens from context $C$ as anchor tokens $S\subseteq C$. We believe that a good selection strategy benefits SAC. Following EPL, our default strategy divides the entire context $C$ into $|S|$ chunks and selects the middle token from each chunk. This setting helps maximize coverage of context $C$. As illustrated in Figure 3a, selected tokens $c_{i}\in S$ are enhanced with anchor embeddings $e_{A}$, yielding an embedding sequence $E=(e_{1},e_{2},\ldots,e_{|C|})$:
$e_{i}=\mathbf{Emb}(c_{i})+\mathbf{1}_{c_{i}\in S}\cdot e_{A}$ (1)
where $\mathbf{1}_{c_{i}\in S}$ is an indicator function that equals 1 when $c_{i}\in S$ and 0 otherwise. Following previous works, we employ a LLM with LoRA parameters $\theta_{LoRA}$ as the compressor: $\tilde{M}=\mathcal{E}(C)=\mathbf{LLM}(E|\theta_{LoRA})$.
While using original tokens from the context avoids learning compressed tokens from scratch and potentially improves the learning efficiency. We notice that because the encoder uses causal attentions, the anchor tokens $S$ do not have visibility to the full sentence, limiting its representation power. Hence we modify the LLM from using causal attention to use bidirectional attention (see Figure 3b), enhancing the LLM’s encoding capability. $\tilde{M}$ can be either the output of anchor tokens from the LLM’s final layer or the Key-Value pairs from each layer. Following 500xCompressor, we use Key-Value pairs as the compressed representation $\tilde{M}$. During pretraining, we only use $\mathcal{L}_{\mathbf{LM}}$ and do not use $\mathcal{L}_{\mathbf{AE}}$ to train the compressor. Following previous work, we use $\mathcal{L}_{\mathbf{QA}}$ for finetuning.
Figure 3: Key differentiators within SAC model architecture. (a) Representative tokens are transformed into anchor tokens through anchor embeddings. (b) The encoder in SAC adopts bidirectional attention, while the decoder operates with causal attention.
4 Experiments
4.1 EXPERIMENTAL SETTING
Dataset. For continued pretraining, we utilize the large-scale corpus  SlimPajama-6B (Soboleva et al., 2023). During finetuning and evaluation, we employ the standard MRQA (Fisch et al., 2019b) question-answering dataset, which consolidates multiple QA tasks and standardizes them into a unified format. We evaluate SAC on both test sets, namely in-domain (ID) and out-of-domain (OOD), to comprehensively assess its in-distribution fitting ability and cross-domain generalization performance.
Implementation Details. SAC utilizes Llama-3.2-1B (Grattafiori & Dubey, 2024) as both the encoder and target LLM. The encoder is equipped with trainable LoRA (Hu et al., 2022) adapters (rank = 128, $\alpha=256$), while the target LLM parameters remain frozen. For each context, we partition it into sub-contexts of 510 tokens each. The compressor compresses each sub-context into a sub-compressed representation, and subsequently concatenates these sub-compressed representations to form the complete compressed representation. The number of anchor tokens $|S|=\lfloor L/r\rfloor$ is determined by the compression ratio $r$ and the length of the sub-context $L$. We train all models in two stages: pretraining for 20,000 optimization steps followed by finetuning for an additional 20,000 steps, both conducted with a batch size of 16. Complete hyperparameter configurations are provided in Appendix A.
Baselines. We use the Llama-3.2-1B model trained on the MRQA (Fisch et al., 2019b) dataset as an uncompressed baseline(denoted as "Full-FT"). We compare our method against several context compression techniques. For hard compression, we choose LLMLingua-2 (Pan et al., 2024b) and evaluate its performance on the Full-FT model. For soft compression, we select ICAE (Ge et al., 2024), 500xCompressor (Li et al., 2025b), and EPL (Zhao et al., 2025). To ensure a fair comparison, all these soft compression baselines are trained on the same dataset as our SAC method.
4.2 Finetuning Results
Tables 1 and 2 report the evaluation results of SAC on in-domain and out-of-domain MRQA datasets, which we analyze from three perspectives: overall performance, effect of compression ratio, and domain generalization.
Overall Performance. SAC consistently outperforms all baselines across a variety of conditions, including compression ratios, and both in-domain and out-of-domain tests, as shown in Tables 1 and 2. Averaging the results of the context compression methods across different compression ratios, SAC shows a maximum improvement of 24.6% F1 / 28.6% EM and a minimum improvement of 4.6% F1 / 5.7% EM in in-domain evaluations. For out-of-domain tests, the maximum improvement is 32.5% F1 / 36.2% EM, with a minimum improvement of 4.6% F1 / 6.9% EM.
Table 1: For the finetuning results, we report in-domain performance using ROUGE-1 F1 (Lin, 2004) and exact match (EM) (Maalouly, 2022) scores on the following datasets: SQuAD (Rajpurkar et al., 2016), NewsQA (Trischler et al., 2017), TriviaQA (Joshi et al., 2017), SearchQA (Dunn et al., 2017), HotpotQA (Yang et al., 2018), and NaturalQuestions (NQ) (Kwiatkowski et al., 2019).
Methods SQuAD NewsQA TriviaQA SearchQA HotpotQA NQ Avg
F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM
Full-FT 77.69 59.71 63.5 46.04 68.80 60.54 73.25 62.07 74.78 59.26 71.01 53.47 71.51 56.85
Lingua-2 32.93 19.57 26.78 13.20 9.67 8.12 45.4 31.80 36.1 22.05 40.08 22.01 31.83 19.46
5x compression constraint
ICAE 36.20 22.12 28.06 13.77 54.63 45.59 65.12 53.06 48.79 33.40 52.36 34.99 47.53 33.82
500x 51.62 33.63 39.70 22.63 57.62 48.76 66.43 54.38 59.10 42.20 57.11 39.26 55.26 40.14
EPL 64.72 44.28 48.74 27.45 63.75 54.54 69.69 57.73 67.16 49.79 63.32 44.16 62.90 46.33
SAC 65.37 44.83 49.39 27.14 65.06 55.93 69.99 58.06 67.41 50.28 64.56 45.44 63.63 46.95
15x compression constraint
ICAE 31.90 18.91 25.25 11.97 51.78 42.94 64.81 52.89 45.22 30.32 48.01 30.67 44.50 31.28
500x 40.68 24.97 32.01 16.76 53.84 44.86 65.65 53.70 53.01 36.30 50.93 33.26 49.35 34.98
EPL 44.58 27.91 33.34 16.69 56.16 47.09 66.36 54.13 54.88 38.38 53.80 35.71 51.52 36.65
SAC 47.43 30.25 36.55 18.07 61.13 52.19 68.97 56.76 58.83 41.86 56.79 38.88 54.95 39.67
51x compression constraint
ICAE 26.17 14.58 22.48 9.69 47.62 39.23 64.31 52.80 38.91 24.78 42.87 26.86 40.39 27.99
500x 30.09 17.11 25.06 12.20 50.84 42.13 64.92 53.29 42.15 27.32 46.07 29.53 43.19 30.26
EPL 30.09 17.49 24.49 11.54 51.15 42.38 65.12 53.16 42.19 27.23 46.29 29.77 43.22 30.26
SAC 31.81 18.78 27.36 13.56 56.73 47.85 65.82 53.76 48.28 32.84 48.22 31.70 46.37 33.08
[…]
Cross-Domain Generalization. We evaluated the generalization capability of SAC on out-of-domain datasets, as shown in Table 2. Under all compression ratio constraints, SAC consistently achieves the highest average F1/EM scores among all methods. Specifically, at a 5x compression ratio, SAC attains average F1 and EM scores of 47.72 and 32.30, outperforming the second-best EPL method by 0.77 and 1.0 points, respectively. At a more challenging 15x compression ratio, SAC achieves average F1 and EM scores of 39.26 and 26.02, surpassing EPL by 2.52 and 2.19 points, with an EM improvement approaching 10%. Even at an extreme 51x compression ratio, SAC maintains average F1 and EM scores of 32.24 and 21.44, still leading EPL by 2.02 and 1.96 points, respectively. These results indicate that the compressed representations learned by SAC exhibit strong cross-domain robustness.
Table 2: For the finetuning results, we report out-of-domain performance using ROUGE-1 F1 and exact match (EM) scores on the following datasets: BioASQ (Tsatsaronis et al., 2015), DROP (Dua et al., 2019), DuoRC (Saha et al., 2018), RACE (Lai et al., 2017), Relation Extraction (RE) (Levy et al., 2017), and TextbookQA (TQA) (Kembhavi et al., 2017).
Methods BioASQ DROP DouRC RACE RE TQA Avg
F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM
Full-FT 49.37 36.77 44.67 34.46 48.82 35.51 35.57 9.64 83.34 72.46 53.32 32.4 52.51 36.87
Lingua-2 27.76 19.48 27.28 18.83 27.07 18.32 17.54 4.15 39.30 20.59 28.42 15.83 27.90 16.20
5x compression constraint
ICAE 36.08 26.06 28.95 21.09 16.67 10.79 15.65 3.12 54.73 41.01 35.24 20.96 31.22 20.51
500x 40.30 28.99 35.40 25.55 29.43 19.32 21.57 4.90 65.43 50.88 38.62 22.75 38.46 25.40
EPL 46.05 32.58 39.94 28.94 39.10 27.12 30.99 6.08 76.07 62.31 49.54 30.74 46.95 31.30
SAC 44.66 31.45 41.55 30.87 39.48 26.92 30.53 6.23 77.87 65.40 52.24 32.93 47.72 32.30
15x compression constraint
ICAE 35.51 24.47 30.39 21.96 13.78 9.06 15.21 3.71 55.24 40.33 34.75 21.56 30.81 20.18
500x 36.30 25.93 33.46 23.55 20.53 12.72 18.49 3.41 54.37 41.11 41.09 25.82 34.04 22.09
EPL 40.52 28.52 32.16 22.29 25.70 16.39 20.97 4.01 59.75 46.34 41.31 25.42 36.74 23.83
SAC 41.31 28.66 36.72 27.48 28.94 18.99 23.35 4.90 61.04 47.90 44.21 28.21 39.26 26.02
51x compression constraint
ICAE 33.82 23.67 27.94 19.29 11.14 6.86 14.89 3.41 47.02 34.02 33.08 19.83 27.98 17.85
500x 32.17 23.07 30.11 21.76 13.42 8.53 15.18 2.67 54.62 41.86 37.10 22.62 30.43 20.09
EPL 32.52 22.21 29.64 20.89 13.16 8.13 17.15 3.12 53.72 40.37 35.15 22.16 30.22 19.48
SAC 36.95 26.86 29.52 20.89 21.85 14.26 15.87 4.00 48.19 36.43 41.05 26.21 32.24 21.44
4.3 Ablation Study
To verify the effectiveness of each key component and strategy in the SAC architecture, we conduct three groups of ablation studies, all performed under a $5\times$ compression ratio.
Component Ablation. As shown in Table 3, our ablation study clearly demonstrates the critical roles of the bidirectional attention and anchor embedding. Removing either component leads to substantial performance degradation in both in-domain (ID) and out-of-domain (OOD) settings. The bidirectional attention mechanism enables anchor tokens to more effectively integrate information from the entire context, producing compressed representations that are more beneficial for downstream tasks. Meanwhile, the anchor embedding provides explicit structural signals that guide the model to accurately identify and process these key tokens, thereby ensuring the effectiveness of information compression.
Table 3: Component ablation results. We report the average F1/EM performance of the model on in-domain (ID) and out-of-domain (OOD) tasks after removing the bidirectional attention (w/o mask) and the anchor embedding (w/o anchor). Full results on all tasks are provided in the Appendix B.2.
Methods ID OOD
TriviaQA HotpotQA Avg BioASQ TextbookQA Avg
F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM
SAC 65.06 55.93 67.41 50.28 66.24 53.11 44.66 31.45 52.24 32.93 48.45 32.19
SAC(w/o mask) 62.60 53.27 64.63 47.43 63.62 50.35 41.93 30.65 48.29 29.67 45.11 30.16
SAC(w/o anchor) 63.90 54.81 65.25 48.31 64.58 51.56 43.70 31.78 51.59 32.20 47.65 31.99
Token Selection. As shown in Table 4, our ablation study investigates the effect of different token selection strategies on the performance of SAC. The results indicate that random selection (Random) significantly degrades performance, not only because the selected tokens lack importance, but also due to their positional randomness, which leads to insufficient global coverage and fails to effectively represent the context. In contrast, both information-based selection (Lingua-2) and our default strategy achieve near-optimal results, and both substantially outperform existing baselines in Tables 1 and 2. This demonstrates that the SAC architecture can effectively leverage and enhance any high-quality token selection strategy, rather than relying on a specific choice, highlighting the generality and robustness of the SAC framework.
Table 4: Token selection ablation results. This table demonstrates how different token selection strategies affect model performance, comparing Random selection, Lingua-2-based selection (Pan et al., 2024b), and our uniform selection (Zhao et al., 2025). We report average F1/EM scores across in-domain (ID) and out-of-domain (OOD) tasks, with comprehensive results for all individual tasks presented in the Appendix B.2.
Methods ID OOD
TriviaQA HotpotQA Avg BioASQ TextbookQA Avg
F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM
SAC 65.06 55.93 67.41 50.28 66.24 53.11 44.66 31.45 52.24 32.93 48.45 32.19
SAC(Random) 59.24 50.22 58.84 41.86 59.04 46.04 43.18 30.59 45.36 29.27 44.27 29.93
SAC(Lingua-2) 64.55 55.13 67.05 49.74 65.80 52.44 44.49 31.91 51.46 32.07 47.98 31.99
AE Effect. As shown in Table 5, we compare the effect of introducing an autoencoding (AE) objective during training on the performance of SAC. Traditional context compression methods employ AE tasks to force independent compression tokens to attend to the original context for reconstruction. However, we find that the AE objective itself has inherent limitations, as its reconstruction target is misaligned with downstream tasks. The experimental results validate this observation: training with only the AE objective leads to a substantial performance drop, and even when combined with the LM objective, the performance still lags behind the full SAC model. This highlights the architectural advantage of SAC: since anchor tokens are naturally semantically aligned with the original context, our method does not require AE objectives to force learning. Instead, SAC effectively aggregates contextual information into anchor token representations solely through anchor embeddings and bidirectional attention. It is worth noting that the ablation experiments in ICAE demonstrate that combining autoencoding tasks with language modeling tasks yields better results (Ge et al., 2024). However, our reproduction on 500xCompressor does not fully support this finding. Specifically, under 15x and 51x compression ratios, 500xCompressor achieves better in-distribution (ID) results when using language modeling tasks alone, with other scenarios being exceptions. This observation raises questions regarding the necessity of autoencoding tasks and suggests that autoencoding may not be entirely essential for context compression methods.
Table 5: Ablation study on the effects of autoencoding (AE) and language modeling (LM) objectives.
Methods ID OOD
5x 15x 51x 5x 15x 51x
F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM
SAC 63.63 46.95 54.95 39.67 46.37 33.08 47.72 32.30 39.26 26.02 32.24 21.44
500x(w/ LM only) 53.23 38.70 49.76 35.71 44.46 31.41 38.22 25.73 33.99 22.05 30.09 18.99
500x(w/ AE+LM) 55.26 40.14 49.35 34.98 43.19 30.26 38.46 25.40 34.04 22.09 30.43 20.09
SAC(w/ AE only) 56.55 40.34 49.93 35.33 43.95 30.64 42.08 27.98 35.50 23.29 28.77 18.32
SAC(w/ AE+LM) 62.04 45.80 51.73 36.67 44.69 31.37 47.26 32.25 37.23 23.96 31.01 19.90
5 ANALYSIS
5.1 Attention Visualization
To understand the unique behavior of compressed models, we analyzed the attention patterns of the final layer at a 5x compression rate. Attention maps for other compression rates can be found in Appendix  C.2.
As observed in Figure 4, The attention map for the 500xCompressor exhibits a distinct anti-diagonal trend. To complete the autoencoding task, the model must condense the entire original sequence into these remaining compressed tokens. This forces later compressed tokens to break locality constraints and actively seek out and attend to distant but important tokens in the sequence. In contrast, the SAC model demonstrates a clear diagonal pattern, where its anchor tokens effectively attend to their neighboring original context tokens, showing a better ability to focus on local information.
Figure 4: Attention maps of different models finetuned under a 5x compression rate. From top to bottom, the figure displays the final layer attention maps for the 500xCompressor and SAC models, respectively. The x-axis represents the original context tokens, and the y-axis represents the compression/anchor tokens.
5.2 Representation Analysis
Key Representation Analysis.In the Key representation space (see Figure 5), the compression tokens (orange) of SAC and EPL are distributed relatively close to the context tokens (blue), while the compression tokens of 500xCompressor are clearly separated from the context tokens. This discrepancy arises from the architectural design of each method. Specifically, although the positional IDs of 500xCompressor’s additional compression tokens are contiguous, their semantics are not aligned, leading to a complete separation in the Key space. In contrast, EPL modifies the positional IDs of its additional compression tokens to share the same rotational angle (RoPE) as the original context tokens, thereby reducing the distance between them. However, in SAC, the anchor tokens are directly embedded within the original context, and their representations maintain close semantic ties with the context tokens from the outset, naturally preventing significant representational divergence.
Value Representation Analysis. In the Value representation space (see Figure 5), the anchor tokens of SAC are uniformly distributed across all regions with the Value representations of the context tokens, without forming independent sub-clusters. This suggests that SAC’s anchor embedding strategy allows for compressed Value representations that more closely match the distribution of the original Value space. In contrast, although EPL’s compression tokens also overlap with the context tokens, their distribution is less complete than SAC’s: they appear relatively sparse in the core regions and show a slight clustering tendency at the boundaries. This indicates that EPL’s Value representations still exhibit a degree of semantic shift relative to the original Value space, which is even more pronounced in the 500xCompressor.
Figure 5: The t-SNE visualization shows the key representations of the final layer KV values for SAC, 500xCompressor (Li et al., 2025b), and EPL (Zhao et al., 2025), respectively.
6 Conclusion
This paper proposes a novel, autoencoding-free context compression method, Semantic-Anchor Compression (SAC), designed to address the performance degradation in downstream tasks caused by context-agnostic compression tokens and autoencoding objectives in existing context compression methods. Unlike traditional context compression approaches, SAC does not rely on training compression tokens to reconstruct the original input. Instead, it directly selects representative anchor tokens from the context and aggregates contextual information into their key-value (KV) representations via a bidirectional attention mechanism. This approach effectively compresses lengthy contexts while avoiding any impairment to the language model’s original language modeling capabilities. Experiments on multiple question answering tasks demonstrate that SAC achieves a high compression ratio and significantly outperforms existing compression methods, highlighting its superiority in balancing compression efficiency and model performance.
7 Reproducibility statement
We declare that the work presented in this paper is reproducible. We provide a link to our anonymous source code as supplementary material: https://anonymous.4open.science/r/SAC-E32C. This code can be used to reproduce the experimental results. The repository includes detailed instructions for environment setup, running experiments, data processing, and result evaluation.
Appendix A Experiment Details
We perform pretraining and finetuning using bf16 precision on 8 NVIDIA RTX 3090 GPUs (24GB). For pretraining, we randomly sample data from the SlimPajama-6B dataset with a token length ranging from 510 to 2040. This data is then split into two halves: one for the auto-encoding (AE) task and the other for the language modeling (LM) task (the AE half is discarded for models without the AE objective). For downstream tasks, we process the MRQA dataset into a (Context, Question, Answer) format for finetuning. Detailed hyperparameters can be found in Table 6.
Table 6: Hyperparameters for training
Hyperparameter Value
Optimizer AdamW
Betas (0.9, 0.95)
Weight decay 0.1
Learning rate 1e-4 (pretrain)
5e-5 (finetuning)
Scheduler Constant
Batch size 16
Warmup 300
Training steps 20k (pretrain)
20k (finetuning)
Clip norm 2.0
Appendix B Detailed Results
B.1 Pretraining Results
As shown in Table 7, our method, SAC, achieves the lowest perplexity (10.79) among all baseline models. This suggests that removing the autoencoding (AE) objective in SAC allows the model to better focus on the language modeling task, thereby improving its predictive capability. Furthermore, since SAC avoids the additional computational overhead from independent compression tokens and the AE task, its training is approximately 31% faster than ICAE and 26% faster than 500xCompressor and EPL.
Table 7: Pretraining comparison of SAC and existing context compression methods, results on LM perplexity and training time.
Methods LM-PPL Training Time(h)
ICAE 12.35 3.85
500xCompress 11.83 3.60
EPL 10.88 3.60
SAC 10.79 2.66
B.2 Ablation Results
In the main text, we have discussed the significant performance gains of SAC over all baseline methods. To provide more detailed evidence, we present the full ablation study results here. As shown in Table 8 and Table 9, our conclusion holds not only in terms of average performance but is also consistently validated on each individual dataset.
Table 8: Ablation studies for SAC under a 5x compression rate on the in-domain dataset are conducted in three sets: component ablation, token selection, and the influence of the auto-encoding (AE) task.
Methods SQuAD NewsQA TriviaQA SearchQA HotpotQA NQ Avg
F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM
Component Ablation
SAC 65.37 44.83 49.39 27.14 65.06 55.93 69.99 58.06 67.41 50.28 64.56 45.44 63.63 46.95
SAC(w/o mask) 60.21 39.93 45.93 25.74 62.60 53.27 66.66 54.82 64.63 47.43 61.53 42.55 60.26 43.96
SAC(w/o anchor) 61.69 41.72 46.52 25.45 63.90 54.81 68.03 56.17 65.25 48.31 62.21 43.88 61.27 45.06
Token Selection
SAC(Random) 52.27 33.41 39.51 19.90 59.24 50.22 68.06 55.87 58.84 41.86 56.57 37.86 55.75 39.85
SAC(Lingua-2) 64.89 44.28 48.92 27.11 64.55 55.13 69.89 58.04 67.05 49.74 64.23 44.93 63.26 46.54
AE Effect
500x(w/ LM only) 44.71 28.89 37.24 20.39 58.97 50.19 65.67 53.74 56.74 40.52 56.07 38.45 53.23 38.70
500x(w/ AE+LM) 51.62 33.63 39.70 22.63 57.62 48.76 66.43 54.38 59.10 42.20 57.11 39.26 55.26 40.14
SAC(w/ AE only) 56.98 37.60 41.09 20.61 58.19 49.08 64.02 51.65 61.58 44.13 57.23 38.98 56.55 40.34
SAC(w/ AE+LM) 64.68 44.62 46.64 25.62 63.34 54.27 68.40 56.48 66.61 49.72 62.56 44.06 62.04 45.80
Table 9: Ablation studies for SAC under a 5x compression rate on the out-of-domain dataset are conducted in three sets: component ablation, token selection, and the influence of the auto-encoding (AE) task.
Methods BioASQ DROP DouRC RACE RE TQA Avg
F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM
Component Ablation
SAC 44.66 31.45 41.55 30.87 39.48 26.92 30.53 6.23 77.87 65.40 52.24 32.93 47.72 32.30
SAC(w/o mask) 41.93 30.65 40.24 28.48 36.48 23.58 28.21 5.49 69.09 55.63 48.29 29.67 44.04 28.92
SAC(w/o anchor) 43.70 31.78 40.55 30.34 36.97 25.58 30.05 6.82 75.88 62.35 51.59 32.20 46.46 31.51
Token Selection
SAC(Random) 43.18 30.59 37.67 27.21 23.44 14.59 22.48 5.64 66.79 51.39 45.36 29.27 39.82 26.45
SAC(Lingua-2) 44.49 31.91 41.50 29.61 39.47 26.58 29.96 7.12 77.67 65.47 51.46 32.07 47.43 32.13
AE Effect
500x(w/ LM only) 43.54 33.11 35.40 25.82 27.71 17.59 19.73 3.86 62.31 48.27 40.60 25.75 38.22 25.73
500x(w/ AE+LM) 40.30 28.99 35.40 25.55 29.43 19.32 21.57 4.90 65.43 50.88 38.62 22.75 38.46 25.40
SAC(w/ AE only) 40.85 29.39 35.32 25.28 31.55 21.32 25.86 4.90 72.29 57.90 46.61 29.08 42.08 27.98
SAC(w/ AE+LM) 44.84 32.31 41.47 31.14 39.29 27.58 30.11 6.23 77.12 64.42 50.74 31.87 47.26 32.26
Appendix C Visualization Analysis
C.1 Training Curves Analysis
Figure 6 shows the training loss curves at different compression ratios on the MRQA dataset. The training loss of our SAC model consistently converges better than other baseline methods across all compression ratios, which demonstrates that the compressed representations obtained from the SAC architecture are more beneficial for language modeling tasks. Notably, as the compression ratio increases appropriately, the difference in convergence between SAC and the other baselines becomes more significant.
Figure 6: Supervised finetuning loss curves. The figure illustrates the training loss trajectories of different models under three compression ratios: 5x, 15x, and 51x.
C.2 Attention Analysis
At a lower 5x compression rate, as shown in Figure 7, the attention map of EPL presents a clear positive diagonal, indicating that its compressed tokens primarily attend to local tokens. In contrast, the attention map of 500xCompressor appears more diffused, while our SAC model exhibits a sparse and highly focused attention pattern, with its anchor tokens attending to only a few key original context tokens. This phenomenon becomes more pronounced with increasing compression rates, being most evident at the 51x compression rate in Figure 9, which strongly demonstrates the robustness of SAC in extreme compression environments.
Figure 7: Attention maps of different models finetuned under a 5x compression rate. From top to bottom, the figure displays the final layer attention maps for the 500xCompressor, EPL, and SAC models, respectively. The x-axis represents the original context tokens, and the y-axis represents the compression tokens.
Figure 8: Attention maps of different models finetuned under a 15x compression rate. From top to bottom, the figure displays the final layer attention maps for the 500xCompressor, EPL, and SAC models, respectively. The x-axis represents the original context tokens, and the y-axis represents the compression/anchor tokens.
Figure 9: Attention maps of different models finetuned under a 51x compression rate. From top to bottom, the figure displays the final layer attention maps for the 500xCompressor, EPL, and SAC models, respectively. The x-axis represents the original context tokens, and the y-axis represents the compression tokens.
Appendix D The Use of Large Language Models
We used a large language model (LLM) as a general-purpose assist tool. The LLM’s primary role was in assisting with writing and text editing, such as refining prose and correcting grammar and spelling to ensure the paper’s professionalism and fluency. We explicitly state that the LLM was not involved in the core ideation or methodological design of this research. All core contributions of the paper, including the proposal of the methodology, the construction and execution of experiments, and the analysis of results, were performed independently by the authors.